# Bridging Entanglement and Magic Resources through Operator Space

### v1

@@ Abstract
For any $N$-qubit unitary $U$, any initial operator $O\in\mathcal{P}_{N}\backslash\{\mathds{1}\}$, for any $\alpha\geq 0$, any bipartition $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{\bar{A}}$, and any unitary propagator $U$, the LOE $\alpha$-Rényi entropy $E^{(\alpha)}_{A}(O_{U})$ satisfies the following inequalities,
$F\underset{(\alpha\leq 2)}{\underset{(U\sim\mu)}{\leq}}E^{(\alpha)}_{A}(O_{U})\leq{M}^{(\alpha)}(O_{U})\leq\nu(U)\leq\tau(U),$ (9)
where the lower bound is valid on average for $\alpha\leq 2$, for sampling over either the $T$-doped Clifford ($\mu_{\tau}$) or $\nu$-compressible ($\mu_{\nu}$) ensembles, and is equal to
$0\leq F:=\begin{cases}\nu-2-\mathcal{O}\left(\frac{2^{\nu}}{D}+\frac{1}{4^{\nu}}\right),&U\sim\mu_{\nu}\\ \log\!\left(\frac{4}{3}\right)\tau-\mathcal{O}\left(\frac{(4/3)^{\tau}}{D}\right),&U\sim\mu_{\tau}.\end{cases}$ (10)
@@
The upper bound is proven through first bounding LOE by OSE, which is apparent from entanglement being the optimum participation entropy over all product bases (with $\mathcal{P}_{N}$ being one such basis). For the remaining two upper-bounds, we use a result from Ref. Jiang and Wang (2023) that $\nu(U)\leq\tau(U)$, and use Prop. 1 to show additionally that OSE lower-bounds the unitary nullity. The lower-bounds on the LOE come directly from Eqs. (3)-(6), after applying Jensen’s inequality for the negative logarithm involved in relating purity to $2-$Rényi entropy, and using the hierarchy of Rényi entropies: $S^{(\alpha)}\leq S^{(\beta)}$ for $\alpha\geq\beta$. The full proof of the upper and lower bounds can be found in Appendices A and B respectively. ∎
Examining both the upper and lower bounds of LOE in Eq. (9), we can see that in the average case for large $N\gg\{\nu,\tau\}\gg 1$, magic resources and operator entanglement approximately coincide. It is instructive to compare Eq. (9) to the equivalent quantities in the Schrödinger picture. Consider the examples of so-called magic states, |$\ket{H}^{\otimes N}:=T^{\otimes N}\ket{++\dots+}$ and random stabilizer states, |$\ket{\psi_{\mathrm{stab}}}:=C\ket{00\dots 0}$ for some deep circuit $C\sim\mathcal{C}_{N}$. In the first case, the aptly named magic state requires very high magic resources to construct (with a $T$-count and state nullity of $\tau(\ket{T}^{\otimes N})=\nu(\ket{T}^{\otimes N})=N$) but has zero entanglement. On the other hand, a random stabilizer state almost surely has near-maximal entanglement, yet requires no magic resources. The existence of the upper bound Eq. (9) in operator space is therefore striking, pointing towards a unified view of non-classical resources in the Heisenberg picture.
To unpack the operational implications of Thm. 2, we return to the physical meaning of LOE. Exclusively for quantum chaotic dynamics, the LOE has been observed to grow extensively with time, while for integrable dynamics it grows at-fastest logarithmically. This conjecture has been confirmed analytically in free-fermion Hamiltonians Prosen and Znidarič (2007); Prosen and Pižorn (2007); Dubail (2017), the Rule-54 interacting-integrable spin chain Alba et al. (2019), and both integrable and chaotic dual unitary circuit models Bertini et al. (2020a, b), with further supporting numerical evidence to be found in Refs. Pižorn and Prosen (2009); Jonay et al. (2018); Alba (2021). Moreover, fast-growing LOE necessarily indicates that the dynamics are scrambling, as quantified by out-of-time-ordered correlators (OTOCs) Dowling et al. (2023). Within the context of the above, we can interpret Thm. 2 via the necessary growth of OSE, nullity, and $T$-count required to simulate chaotic dynamics.
@@
Note that due to its inherent Lieb-Robinson light-cone for the OSE magic monotone (forming the tightest upper bound in Eq. (9)), $\mathcal{O}(t)$ scaling is maximal for local Hamiltonian dynamics Dowling et al. (2024). It is worth comparing the above corollary to the results of Refs. Haferkamp et al. (2022a); Leone et al. (2021). There, it was found that for the $T$-doped Clifford ensemble, $\mathcal{O}(N)$ $T$-gates are both necessary Haferkamp et al. (2022a) and sufficient  Leone et al. (2021) to generate an $\epsilon-$approximate $t$-design, for a sufficiently small error in order to differentiate such randomness using e.g. $8$-point OTOCs. However, these results are valid on-average over unitary ensembles, and so leave open the question of what to expect for deterministic dynamics, such as those generated by some non-integrable spin-chain Hamiltonian. Corollary 3 therefore complements these results to assure us that in fact, quantum chaos really cannot be simulated classically.
To elucidate the preceding statement, from Thm. 2 we can also immediately deduce a hierarchy of computational complexities of simulation using (apparently) inequivalent methods. We consider three prominent techniques in many-body dynamical simulation: the tensor network method of Heisenberg picture time-evolving block decimation (H-TEBD) Hartmann et al. (2009) has a computational cost that scales (exponentially) with LOE Verstraete and Cirac (2006); Schuch et al. (2008); stabilizer methods scale (exponentially) with $T$-count according to the Gottesmann-Knill theorem Gottesman (1998); Aaronson and Gottesman (2004); and Pauli truncation methods which rely on the biased suppression of high-weight Pauli terms in locally noisy/scrambling Heisenberg evolution Rakovszky et al. (2022); Lloyd et al. (2023); Begusic et al. (2024); Srivatsa et al. (2024); Begušić and Chan (2024), and have an expense (exponentially) bounded by the OSE Dowling et al. (2024). Note that these resource costs being large do not preclude the efficient simulation of certain features of a system to polynomial precision, such as the anti-concentration of local observables being well-approximated by either: Pauli-truncating an operator to low OSE Angrisani et al. (2024), or analogously using area-law entangled random tensor-networks Cheng et al. (2024). In the following, if the resource cost $R$ of a simulation method for some one-parameter unitary $U_{t}$ scales extensively with time [layers] $t$, $R(U_{t})\sim\mathcal{O}(t)$, we say that the dynamics [circuit] is non-simulable with respect to the said method.
@@
If $U_{t}$ is non-simulable according to H-TEBD, then it is also necessarily non-simulable using Pauli truncation, then it is also necessarily non-simulable using stabilizer methods.
@@
The above result means that if a unitary $U_{t}$ is efficiently simulable according to either stabilizer or Pauli truncation methods, then it must also be efficient to model using H-TEBD.
@@
Conclusion.— The local-operator entanglement of a unitary propagator $U$ has long been conjectured as a faithful measure of quantum chaos, alongside its direct interpretation as tensor network simulability of the dynamics. Here, we have shown through Thm. 2 that this quantity is limited by the minimum of $U$’s entangling capacity and its magic. The implications here are two-fold: we concretely bridge two seemingly discordant resources of fundamental importance in the one quantity. But by extension, our results also integrate a new viewpoint on the non-simulability of complex quantum dynamics.
@@
We emphasize, however, that this work does not rule out the possibility of classically simulating volume-law LOE through other means. For example, matchgate circuits constitute an alternate class of classically tractable systems. Here, Gaussian observables of free-fermionic systems can be computed efficiently  Jozsa and Miyake (2008); Reardon-Smith et al. (2024). Moreover, the relevant class has been observed to display both high entanglement and magic Collura et al. (2024). Nevertheless, we conjecture that a similar relation to the upper bound of Thm 2 will hold in this setting: that LOE can be bounded by the non-Gaussian resources Zhuang et al. (2018) employed. This intuition is bolstered by the observation that LOE grows at fastest logarithmically with time for free-fermionic Hamiltonian evolution Prosen and Znidarič (2007); Prosen and Pižorn (2007); Dubail (2017).
@@
Beyond strict limits, it is important to explore further the behavior of LOE in concrete settings. One way to understand the average-case saturation in Thm. 2 is that typical dynamics are maximally entangling. This makes magic the bottleneck, and thus essentially equivalent to the LOE. Less typical scenarios should also be studied, where the relative amounts of each is on equal footing. Appropriate methods on this front could be found in stabilizer tensor network Masot-Llima and Garcia-Saez (2024); Nakhl et al. (2024) and Clifford-assisted matrix-product state Fux et al. (2024b); Qian et al. (2024) techniques. These combine both stabilizer and tensor network principles to more efficiently simulate certain systems with an intermediate amount of magic resource and may be well-suited to studying LOE in practice. Already, interesting regimes have been numerically identified, such as in Refs. Fux et al. (2024b); Nakhl et al. (2024) it was observed that $T$-doped Clifford circuits can be efficiently simulated up to $\tau\lesssim N$. One might like to investigate more deeply the regimes between the points $m\approx\{0,N,2N\}$ for different magic monotones $m$ and varied entanglement. This is particularly interesting in the context of unitary nullity; c.f. Prop. 1 and the operational power of $\nu$-compressible states studied in Ref. Gu et al. (2024). Understanding such restricted-complexity dynamical systems will offer further insight into quantum randomness Haferkamp et al. (2022a); Leone et al. (2021), simulability Fux et al. (2024b); Nakhl et al. (2024), learnability Leone et al. (2023, 2024b); Gu et al. (2024), and integrability Turkeshi et al. (2024); Tirrito et al. (2024b); Odavic et al. (2024b); Dowling et al. (2024).
@@ Acknowledgements.
We are grateful to Lorenzo Leone and Xhek Turkeshi for useful discussions and comments on the manuscript. ND acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769. GALW is supported by an Alexander von Humboldt Foundation research fellowship.
@@ Appendix A Upper-bounding LOE by magic monotones
In this Appendix, we will prove the upper bound of the main result, Thm. 2. This upper-bound is the cumulation of several theorems which we will supply here. We will first prove the nullity compression result of Prop. 1 and discuss its implications in some more detail. The strategy is then to upper-bound the OSE by the unitary nullity, use that nullity lower-bounds the $T$-count, before proving that OSE further always lower-bounds the LOE of the same Rényi index.
A.1 $\nu$-compressibility of unitaries
First, we restate and prove Prop. 1, on the efficient representation of unitaries with restricted magic resource of unitary nullity. Note that this theorem is similar to that given in Refs. Leone et al. (2024a, b); Gu et al. (2024). See 1
Proof.
A unitary $U$ stabilizes $2^{2N-\nu}$ Pauli strings, mapping them to Paulis. Call this subgroup of Pauli strings $\mathrm{Stab}(U)$. There exists a Clifford $C$ which also does this, such that $\forall P\in\mathrm{Stab}(U)$, $C^{\dagger}PC=U^{\dagger}PU$. Now, for some $V\in\mathcal{U}_{N}$, we can write the unitary as $U=CV$, where $V$ necessarily commutes with all elements of $\mathrm{Stab}(U)$. Therefore, $V$ acts non-trivially only on a space spanned by $2^{\nu}$ Pauli strings, mapping to a space also of dimension $2^{\nu}$ by linearity. Then we can choose pre- and post-processing Cliffords $C_{0}$ and $C_{1}$ which map these input and output spaces respectively to spaces spanned only by local Paulis on the first $\nu/2$ qubits. The factor of $1/2$ comes from the fact that each site has $4$ possible local Paulis (or two independent generators, one corresponding to the input space, and one to the output space). We can freely choose $C_{0}$ and $C_{1}$ to act trivially on $\mathrm{Stab}(U)$. Then together with $C$ above, we arrive at the representation Eq. (5). ∎
The above then directly leads to a well-defined ensemble of unitaries, when randomly sampling $C_{0},C_{1}$ over the Clifford group, and $V$ over the unitary group on a restricted space of $\nu/2$ qubits. When taking replica averages according to $\mu_{\nu}$, the exact location of the $\nu$ qubits that $V$ acts on does not matter, as both the unitary and Clifford groups are invariant under left/right multiplication by permutation operations. The decomposition Prop. 1 and the ensemble $\mu_{\nu}$ are of independent interest, and should find application in other areas of many-body dynamics. As mentioned in the main text, the above improves upon the previous unitary compressibility statements of Refs. Leone et al. (2024a, b); Gu et al. (2024) by a factor of $1/2$. It would be interesting to study Prop. 1 in the context of learnability and gate-complexity, e.g. using similar approaches to Refs. Leone et al. (2024a, b); Gu et al. (2024). For instance, we immediately arrive at an upper-bound on the circuit complexity of any unitary with nullity $\nu(U)$. The circuit complexity of a unitary $g(U)$ is defined as the minimum number of elementary gates required to synthesize a unitary, from some universal gate set Haferkamp et al. (2022b).
Corollary 5.
The circuit complexity of an $N-$qubit $U$ with a unitary nullity of $\nu(U)$, is
$g(U)=\mathcal{O}(N^{2}+\exp(\nu/2))$ (11)
Proof.
The above follows immediately from Prop. 1, together with the fact that Clifford unitaries can be compressed to $O(N^{2})$ gates Bravyi and Maslov (2020), whereas arbitrary unitaries can have up to exponential circuit complexity with respect to their support. ∎
Clearly, if the nullity scales logarithmically with $N$, then the unitary necessarily has low (polynomial) circuit complexity. Recalling the bound $\nu(U)\leq\tau(U)$ Jiang and Wang (2023), said another way: circuits with a logarithmic number of non-Clifford gates have a complexity of $\mathcal{O}(N^{2})$, which improves upon the naive scaling of $\mathcal{O}(\log(N)N^{2})$ (which can be deduced from a $T-$doped Clifford circuit representation). Notice that $\mathcal{O}(N^{2})$ scaling is exactly that of Clifford circuits, meaning that a logarithmic number of non-Cliffords do not change the scaling of circuit complexity (albeit the prefators will change).
A.2 Proof of upper bounds
From Prop. 1, we can immediately relate unitary nullity to the OSE.
@@ A.2 Proof of upper bounds
From Prop. 1, we write the unitary propagator as $U=C_{0}VC_{1}$, where $C\in\mathcal{C}_{N}$, and $V$ is an arbitrary unitary acting on at-most $\lceil\nu(U)/2\rceil\leq N$ qubits. Take $O\in\mathcal{P}_{N}$. Then,
$U^{\dagger}OU=C_{1}V^{\dagger}C_{0}^{\dagger}OC_{0}VC_{1}=C_{1}V^{\dagger}O^{\prime}VC_{1}$ (13)
where also $O^{\prime}\in\mathcal{P}_{N}$ by the definition of Clifford unitaries. In the worst case (with respect to maximizing the OSE), the action of $V$ on a Pauli string can output a uniform superposition of at most $4^{\nu/2}$ Pauli strings, as $V$ acts on only $\nu/2$ qubits. The final Clifford conjugation with respect to $C_{1}$ cannot increase the OSE further, as each Pauli string is again mapped to another Pauli. The entropy of a uniform mixture is simply
$\mathcal{M}^{(\alpha)}\left(\frac{1}{2^{\nu}}\sum^{2^{\nu}}_{i=1}P_{i}\right)$ $=\log(2^{\nu})=\nu.$ (14)
From this optimal case, the bound of Eq. (12) follows directly. ∎
@@
Now that we have that OSE is always the smallest of the three magic monotones we use in this work (see Table 1), it remains to prove the relation between OSE and LOE.
Lemma 7.
@@
$M^{(\alpha)}(O_{U})=\frac{1}{1-\alpha}\log\sum_{P\in\mathcal{P}}\left(D^{-1}\operatorname{tr}[O_{U}P]\right)^{2\alpha}=\frac{1}{1-\alpha}\sum_{P\in\mathcal{P}}|\braket{P}{O_{U}}|^{2\alpha}=\frac{1}{1-\alpha}\sum_{i=1}^{4^{N}}|\braket{i}{O_{U}}|^{2\alpha}$ (16)
@@
$|O_{U}\rangle\!\rangle=\sum_{j}\lambda_{j}\ket{a_{j}}\ket{b_{j}}.$ (17)
@@
$|O_{U}\rangle\!\rangle=\sum_{i,j}\lambda_{i,j}\ket{i}\ket{j},$ (19)
and the corresponding entropy $S^{(\alpha)}_{\mathcal{P}}(\ket{O_{t}})$ is the OSE [Eq.(16)]. The Schmidt basis is defined as the unique basis $\mathcal{B}$ satisfying the bipartite-product condition which minimizes the corresponding bipartite-basis entropy. Then the bipartite-basis entropy of the Schmidt basis is just the entanglement entropy, and so lower bounds all other bipartite-basis entropies, including the OSE. ∎
Combining Prop. 6, the fact that $\nu(U)\leq\tau(U)$, and that Lemma 7, we arrive at the hierarchy of upper bounds from Thm. 2,
@@
Appendix B Average-case lower bound on LOE from unitary ensembles
In this section, we will detail the proofs for the lower bounds of average LOE in terms of magic monotones of nullity and $T$-count, over both the $\nu$-compressible and $T$-doped ensembles.

### v2

@@ Abstract
For any $N$-qubit unitary $U$, any initial operator $O\in\mathcal{P}_{N}\backslash\{\mathds{1}\}$, for any $\alpha\geq 0$, any bipartition $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{\bar{A}}$, and any unitary propagator $U$, the LOE $\alpha$-Rényi entropy satisfies,
$E^{(\alpha)}_{A}(O_{U})\leq{M}^{(\alpha)}(O_{U})\leq\nu(U)\leq\tau(U).$ (7)
@@
The upper bound is proven through first bounding LOE by OSE, which is apparent from entanglement being the optimum participation entropy over all product bases (with $\mathcal{P}_{N}$ being one such basis). For the remaining two upper-bounds, we use a result from Ref. Jiang and Wang (2023) that $\nu(U)\leq\tau(U)$, and use a counting argument from the definition of nullity to show that ${M}^{(\alpha)}(O_{U})\leq\nu(U)$. See App. A for details. ∎
Examining both the upper and lower bounds of LOE in Thms. 2 and 1, we can see that in the average case for large $N\gg\{\nu,\tau\}\gg 1$, magic resources and operator entanglement approximately coincide. It is instructive to compare Eq. (7) to the equivalent quantities in the Schrödinger picture. Consider the examples of so-called magic states, |$\ket{H}^{\otimes N}:=T^{\otimes N}\ket{++\dots+}$ and random stabilizer states, |$\ket{\psi_{\mathrm{stab}}}:=C\ket{00\dots 0}$ for some deep circuit $C\sim\mathcal{C}_{N}$. In the first case, the aptly named magic state requires very high magic resources to construct (with a $T$-count and state nullity of $\tau(\ket{T}^{\otimes N})=\nu(\ket{T}^{\otimes N})=N$) but has zero entanglement. On the other hand, a random stabilizer state almost surely has near-maximal entanglement, yet requires no magic resources. The existence of the upper bound Eq. (7) in operator space is therefore striking, pointing towards a unified view of non-classical resources in the Heisenberg picture.
To unpack the operational implications of Thm. 2, we return to the physical meaning of LOE. Exclusively for chaotic dynamics, the LOE has been observed to grow extensively with time, while for integrable dynamics it grows at-fastest logarithmically. This conjecture has been confirmed analytically in free Prosen and Znidarič (2007); Prosen and Pižorn (2007); Dubail (2017), interacting-integrable Alba et al. (2019); Bertini et al. (2020b), and even chaotic Bertini et al. (2020a) spin chains, with further supporting numerical evidence to be found in Refs. Pižorn and Prosen (2009); Jonay et al. (2018); Alba (2021). Moreover, fast-growing LOE necessarily indicates that the dynamics are scrambling, as quantified by out-of-time-ordered correlators (OTOCs) Dowling et al. (2023). Note that LOE should not be confused with the operator entanglement of $U$ Zanardi (2001), which typically grows fast irrespective of integrability Zhou and Luitz (2017); Dubail (2017). Within the context of the above, we can interpret Thm. 2 via the necessary growth of OSE, nullity, and $T$-count under chaotic evolution.
@@
Note that due to its inherent Lieb-Robinson light-cone for the OSE magic monotone (forming the tightest upper bound in Eq. (7)), $\mathcal{O}(t)$ scaling is maximal for local Hamiltonian dynamics Dowling et al. (2024). It is worth comparing the above corollary to the results of Refs. Haferkamp et al. (2022a); Leone et al. (2021). There, it was found that for the $T$-doped Clifford ensemble, $\mathcal{O}(N)$ $T$-gates are both necessary Haferkamp et al. (2022a) and sufficient Leone et al. (2021) to reproduce Haar predicted values for certain measures of chaos, e.g. higher-order OTOCs. However, these results are valid on-average over unitary ensembles, and so leave open the question of what to expect for deterministic (Hamiltonian) dynamics. Corollary 3 therefore complements these results: assuring us that in fact, quantum chaos really cannot be simulated classically.
To elucidate the preceding statement, from Thm. 2 we can immediately deduce a hierarchy of computational complexities using inequivalent methods. We consider three prominent techniques in many-body dynamical simulation: the tensor network method of Heisenberg picture time-evolving block decimation (H-TEBD) Hartmann et al. (2009) has a computational cost that scales (exponentially) with LOE Verstraete and Cirac (2006); Schuch et al. (2008); stabilizer methods scale (exponentially) with $T$-count according to the Gottesmann-Knill theorem Gottesman (1998); Aaronson and Gottesman (2004); Bravyi et al. (2019); Pashayan et al. (2022); and Pauli truncation methods Rakovszky et al. (2022); Lloyd et al. (2023); Begusic et al. (2024); Srivatsa et al. (2024); Begušić and Chan (2024) have an expense bounded (exponentially) by the OSE Dowling et al. (2024). Note that these resource costs being large do not preclude the efficient simulation of certain features of a system to polynomial precision, such as the anti-concentration of local observables being well-approximated by either: Pauli-truncating an operator to low OSE Angrisani et al. (2024), or analogously using area-law entangled random tensor-networks Cheng et al. (2024). In the following, if the (exponential) resource cost $R$ of a simulation method for some one-parameter unitary $U_{t}$ scales extensively with time [layers] $t$, $R(U_{t})\sim\mathcal{O}(t)$, we say that the dynamics [circuit] is non-simulable with respect to the said method.
@@
If $U_{t}$ is non-simulable according to H-TEBD, then it is also necessarily non-simulable using Pauli truncation, and it is further necessarily non-simulable using stabilizer methods.
@@
The contrapositive of the above statement means that if a unitary $U_{t}$ is efficiently simulable according to either stabilizer or Pauli truncation methods, then it must also be efficient to model using H-TEBD.
@@
Conclusion.— The local-operator entanglement of the Heisenberg operator $O_{U}$ has long been conjectured as a faithful measure of quantum chaos of $U$, alongside its direct interpretation as tensor network simulability of the dynamics. Here, we have shown through Thm. 2 that this quantity is limited by the minimum of $U$’s entangling capacity and its magic. The implications here are two-fold: we concretely bridge two seemingly discordant resources of fundamental importance in the one quantity. But by extension, our results also integrate a new viewpoint on the non-simulability of complex quantum dynamics.
@@
We emphasize, however, that this work does not rule out the possibility of classically simulating volume-law LOE through other means. For example, matchgate circuits constitute an alternate class of classically tractable systems. Here, Gaussian observables of free-fermionic systems can be computed efficiently  Jozsa and Miyake (2008); Reardon-Smith et al. (2024). Moreover, the relevant class has been observed to display both high entanglement and magic Collura et al. (2024). Nevertheless, we conjecture that a similar relation to the upper bound of Thm 2 will hold in this setting: that LOE can be bounded by the non-Gaussian resources Zhuang et al. (2018); Bu et al. (2024) employed. This intuition is bolstered by the observation that LOE grows at fastest logarithmically with time for free-fermionic Hamiltonian evolution Prosen and Znidarič (2007); Prosen and Pižorn (2007); Dubail (2017).
@@
Beyond strict limits, it is important to explore further the behavior of LOE in concrete settings. One way to understand the average-case saturation in Thm. 2 is that typical dynamics are maximally entangling. This makes magic the bottleneck, and thus essentially equivalent to the LOE. Less typical scenarios should also be studied, where the relative amounts of each is on equal footing. Appropriate methods on this front include those combining stabilizer and tensor network principles Masot-Llima and Garcia-Saez (2024); Fux et al. (2024b); Qian et al. (2024); Nakhl et al. (2024), suitable for certain systems with an intermediate amount of magic resource. This setting may be well-suited to studying LOE in practice, and already interesting regimes have been numerically identified, such as the efficient simulation of $T$-doped Clifford circuits, where $\tau\lesssim N$ Fux et al. (2024b); Nakhl et al. (2024). Such intermediate-amount of magic is also particularly interesting in the context of unitary nullity; c.f. Eq. (3) and the operational power of $\nu$-compressible states studied in Ref. Gu et al. (2024). Insight on this front may be gleaned from determining classes of unitaries which saturate the respective bounds on compressibility described below Eq. (3). Understanding such restricted-complexity dynamical systems will offer further insight into quantum randomness Haferkamp et al. (2022a); Leone et al. (2021), simulability Fux et al. (2024b); Nakhl et al. (2024), learnability Leone et al. (2023, 2024b); Gu et al. (2024), and integrability Turkeshi et al. (2024); Tirrito et al. (2024b); Odavic et al. (2024b); Dowling et al. (2024).
@@ Acknowledgements.
We are grateful to Lorenzo Leone, Xhek Turkeshi, and Pavel Kos for useful discussions and comments on the manuscript. ND acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769. GALW is supported by an Alexander von Humboldt Foundation research fellowship.
Supplemental Material
@@ Appendix A Upper-bounding LOE by magic monotones
We will here prove the upper bounds of Thm. 2. The strategy is to upper-bound the OSE by the unitary nullity, then use that nullity lower-bounds the $T$-count, before proving that OSE further always lower-bounds the LOE of the same Rényi index.
We recall the definition of unitary nullity for a given $U$, namely: a unitary $U$ stabilizes $2^{2N-\nu}$ Pauli strings, mapping them to Paulis. This subgroup of Pauli strings is called the unitary stabilizer group of $U$, $\mathrm{Stab}(U)$. Using these definitions, we can immediately relate unitary nullity to the OSE.
@@
Taking $O\in\mathcal{P}_{N}$, we consider the conjugate action of $U$ $O_{U}=U^{\dagger}OU$. To derive an upper bound, we consider the ‘worst case’ of $O\in\mathcal{P}_{N}\backslash\mathrm{Stab}(U)$. Otherwise, $O_{U}\in\mathcal{P}_{N}$ and OSE is trivially zero. In the non-trivial case, the output $O_{U}$ is a superposition of at most $2^{\nu}$ terms (the cardinality of $\mathcal{P}_{N}\backslash\mathrm{Stab}(U)$). The highest entropy of the square coefficients of a superposition of $2^{\nu}$ terms is ${M}^{(\alpha)}(O_{U})=\log(2^{\nu})=\nu$, corresponding to a uniform superposition. From this optimal case, the bound of Eq. (8) follows directly. ∎
@@
Recalling that $\nu(U)\leq\tau(U)$ Jiang and Wang (2023), we have that OSE is always the smallest of the three magic monotones studied in this work (see Table 1), it remains to prove the relation between OSE and LOE.
Proposition 6.
@@
$M^{(\alpha)}(O_{U})$ $=\frac{1}{1-\alpha}\log\sum_{P\in\mathcal{P}}\left(D^{-1}\operatorname{tr}[O_{U}P]\right)^{2\alpha}$ (10)
$=\frac{1}{1-\alpha}\log\sum_{P\in\mathcal{P}}|\langle\!\langle{P|O_{U}}\rangle\!\rangle|^{2\alpha}$ (11)
$=\frac{1}{1-\alpha}\log\sum_{i=1}^{4^{N}}|\langle\!\langle{i|O_{U}}\rangle\!\rangle|^{2\alpha}$ (12)
@@
$|O_{U}\rangle\!\rangle=\sum_{j}\lambda_{j}|a_{j}\rangle\!\rangle|b_{j}\rangle\!\rangle.$ (13)
@@
$|O_{U}\rangle\!\rangle=\sum_{i,j}\lambda_{i,j}|i\rangle\!\rangle|j\rangle\!\rangle,$ (15)
and the corresponding entropy $S^{(\alpha)}_{\mathcal{P}}(|{O_{U}}\rangle\!\rangle)$ is the OSE [Eq.(12)]. The Schmidt basis is defined as the unique basis $\mathcal{B}$ satisfying the bipartite-product condition which minimizes the corresponding bipartite-basis entropy. Then the bipartite-basis entropy of the Schmidt basis is just the entanglement entropy, and so lower bounds all other bipartite-basis entropies, including the OSE. ∎
Combining Props. 5 and 6, together with the fact that $\nu(U)\leq\tau(U)$ Jiang and Wang (2023), we arrive at the hierarchy of upper bounds of Thm. 2,
@@
Appendix B Bounding Rényi LOE from average operator purity in replica space
In this document, we will detail the proofs for the lower bounds of average LOE in terms of magic monotones of nullity and $T$-count, over both the $\nu$-compressible and $T$-doped ensembles.